[Hallc_running] [New Logentry] RC update for Hall C, Saturday swing to Sunday swing

mack at jlab.org mack at jlab.org
Sun Mar 8 19:10:01 EDT 2020

Logentry Text:
     Saturday SWING was longitudinal target production at nominal HMS/SHMS -3.5/-3.4 GeV data with 3He target.

    Also Saturday SWING the accelerator had trouble delivering beam at the nominal RF load, so beam currents in Halls A and C were tweaked down. Looks like they went with "Door Number -5%" power reduction as default, but they've been pushing currents higher when the cavities are behaving. https://logbooks.jlab.org/entry/3799558 https://logbooks.jlab.org/entry/3799566 

    Our beloved Carlos G on OWL keeps noting confusion in the CODA run stamp regarding the half wave plate state. https://logbooks.jlab.org/entry/3799706 Please skim the hclog when you come in and show him we not only care but are "COO page 8 section 4.6"-compliant : " read the logbook to be aware of changes in goals, operating parameters, and new documentation".    :) 

    Target was switched to transverse early on Sunday day shift. (This takes more time now since an interlock has to be moved. https://logbooks.jlab.org/entry/3799915 ) After putting the SHMS quads on loop (see below), steady production HMS/SHMS -3.5/-3.4 GeV data with 3He target. Xiaochao implies we are transverse-hungry, so we will be spending roughly 20 hours on this transverse setting. 

    When Xiaochao came on shift Sunday day shift, her tingling spider senses told her that the SHMS quads had not been cycled yesterday, and this was eventually confirmed with strip charts. (Maybe we can ask Mike Fowler for a "cycled" flag on the magnet screens for Xmas?) According to Holly's SHMS opus at https://hallcweb.jlab.org/doc-private/ShowDocument?docid=1007 , cycling is a concern for Q1 since it derives a lot of field from iron. Fortunately, and amazingly, the Q1 Hall probe fields were the same to 1G out of 5036 G. https://logbooks.jlab.org/entry/3799898 But to make sure incoming data were clean, SHMS quads were cycled and reset. https://logbooks.jlab.org/entry/3799907 (Of course, even if the field is the same, we cycle magnets to reproduce the effective length as well.) 

    Four target polarization measurements since restoration yesterday: 52.7%(L), 53.5%(L), 53.9%(L), and 51.9%(T). 

    A propos of scraping discussions the other day, Simona posted some nice FR plots on mixed targets : https://logbooks.jlab.org/entry/3800116


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3800145
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200308/3173b8d2/attachment.html>

More information about the Hallc_running mailing list